# Entropy-Based Decoding for Retrieval-Augmented Large Language Models

### v1

@@ Abstract
Augmenting Large Language Models (LLMs) with retrieved external knowledge has proven effective for improving the factual accuracy of generated responses. Despite their success, retrieval-augmented LLMs still face the distractibility issue, where the generated responses are negatively influenced by noise from both external and internal knowledge sources. In this paper, we introduce a novel, training-free decoding method guided by entropy considerations to mitigate this issue. Our approach utilizes entropy-based document-parallel ensemble decoding to prioritize low-entropy distributions from retrieved documents, thereby enhancing the extraction of relevant information of context. Additionally, it incorporates a contrastive decoding mechanism that contrasts the obtained low-entropy ensemble distribution with the high-entropy distribution derived from the model’s internal knowledge across layers, which ensures a greater emphasis on reliable external information. Extensive experiments on open-domain question answering datasets demonstrate the superiority of our method.
@@
Zexuan Qiu1 Zijing Ou2 Bin Wu3 Jingjing Li1 Aiwei Liu4 Irwin King1 1The Chinese University of Hong Kong 2Imperial College London 3University College London 4Tsinghua University
@@ 1 Introduction
How to eliminate the impact of the above-mentioned distractibility issue, so as to extract useful knowledge from the retrieved context for the input query, is our research focus. Although existing works strive to effectively leverage the retrieved context by directly fine-tuning retrieval-augmented LLMs (Lin et al., 2023; Zhang et al., 2024) or incorporating trainable encoder modules (Izacard and Grave, 2020; Yen et al., 2024), these approaches require additional training, rendering them potentially impractical in resource-constrained environments. In this paper, we propose a novel decoding method guided by entropy considerations to simultaneously mitigate the impact of noisy information from both the external context and parametric knowledge. The proposed method can be seamlessly integrated into LLMs without requiring additional tuning.
@@ 2.1 Entropy-Based Document Ensemble
There are multiple choices to compute the weights. A straightforward option is to use uniform weighting, i.e., $w_{j,t}=\frac{1}{|D|}$. However, this method may fail to effectively extract valuable information when irrelevant documents are included among the top-$K$ retrieved documents. Another option is to utilize a similarity score $s(d_{j},x)$ between the query and the retrieved document, e.g., BM25, as the time-independent ensemble weights $w_{j,t}\propto s(d_{j},x)$ (Lewis et al., 2020). However, this retriever-based scoring apporach may hinder the LLM’s ability to extract relevant information when the retrieved documents contain numerous distractors, drastically reducing the factual accuracy of the responses as illustrated in § 4.1.
@@ 3 Related Works
The idea of contrastive decoding (CD) has been previously applied in controllable text generation to produce non-toxic by DExperts (Liu et al., 2021). Later, Li et al. (2022) formalized CD as a method to enhance open-ended text generation without any additional training by maximizing the difference in log probabilities between an expert LLM and an amateur LLM. This approach has demonstrated strong performance in various domains, including reasoning (O’Brien and Lewis, 2023) and neural machine translation (Waldendorf et al., 2024). CD can also be interpreted as maximizing pointwise mutual information (PMI), which has proven effective in other contexts. For instance, Li et al. (2016) use a training objective that maximizes PMI to generate more diverse conversational responses, while CAD (Shi et al., 2023b) employs a PMI-adjust output probability distribution to resolve the knowledge conflict. Chuang et al. (2023) propose a decoding strategy that contrasts different layers of the same LLM to more effectively highlight factual knowledge obtained during pre-training. Similar principles are also applied in visual LLMs, where Leng et al. (2023) mitigate object hallucination by contrasting distributions derived from original and distorted visual inputs. Alternatively, our proposed CLeHe leverages layer-wise entropy-based contrastive decoding to prioritize external knowledge over the parametric knowledge inherent in the LLM itself.
@@ 4 Experiments
We compare three training-free baselines. Naive: this method concatenates all retrieved documents directly along with the question to form the prompt for LLMs; RePlug (Shi et al., 2023c): it utilizes a normalized retriever weight to ensemble during the decoding process; AvgEns : it follows the formulation in Eq.(2) and assigns the same weight to each document during each generation step. We verify the effectiveness of these decoding methods on 4 LLMs: Llama-2-7B, Llama-2-13B, Mistral-7B-v0.1, and Llama3-8B. Among them, Llama-3-8B supports an 8K context length, while the other three LLMs support a 4K context length. Few evaluation samples exceed 4k in context length, and they are removed to ensure the validity of the assessment.
@@
Our method introduces two hyperparameters: $\tau$ to control the relative importance of different documents during decoding; and $\beta$ balancing contextual and parametric knowledge. We extract a subset from the WebQ training set for validation to determine the hyperparameter value for each LLM. Ultimately, $\tau$ is set as 0.25 for Llama-3-8B and as 0.1 for the other three models. For $\beta$, 5.0 is chosen for Llama-2-7B, while the other models are assigned a value of 0.25. During the inference, greedy decoding is utilized for reproducibility. When looking for the layer with the highest entropy, we focus our search exclusively on the candidate layers. In our preliminary experiments, this approach improves computational efficiency and slightly enhances model performance. For Llama-2-7B, Mistral-7B-v0.1, and Llama-3-8B with 32 hidden layers, the candidate layers are set to $\{17,\dots,32\}$, and only even-numbered layers will be searched. For Llama2-13B with 40 hidden layers, the candidate layers are set to $\{31,\dots,40\}$. All experiments are conducted on a single A100 80GB GPU.
@@ 4.1 Analyzing the Distraction Phenomenon in Retrieved Context
We are particularly interested in a challenging QA scenario proposed by  (Liu et al., 2024), in which the oracle document is surrounded by numerous semantically similar distractor documents. This configuration challenges the efficacy of the Naive LLM-RAG, preventing it from accurately deriving answers based on contextual information, thereby leading to the “Lost in the middle” phenomenon. Specifically, following (Liu et al., 2024), given a query from NaturalQuestions-Open (Kwiatkowski et al., 2019), we select a Wikipedia paragraph containing the answer from the NaturalQuestions annotations as the oracle document. Then, Contriever (Gautier et al., 2022) is employed to extract K-1 additional paragraphs from the Wikipedia corpus that are highly relevant to the query yet do not include the ground truth answer, functioning as distractor documents. The query, the oracle document, and K-1 distractor documents are subsequently processed by the LLM to generate an answer.
From Figure 2, it is evident that the performance of the Naive method which concatenates all documents to compose the context, is highly sensitive to the placement of the oracle document within the retrieved documents. Among the four evaluated LLMs, the performance of the Naive method significantly deteriorates when the oracle document is neither at the very beginning nor at the end. In contrast, since the proposed LeEns  processes each document in parallel during decoding, its performance is naturally independent of the position of the oracle document. In almost all positions, LeEns substantially surpasses the performance of the Naive method. Notably, in this challenging scenario, RePlug which ensemble documents’ distributions based on retriever weights perform exceedingly poorly, achieving results merely on par with AvgEns . Based on this observation, we further conduct the weight analysis in Figure 3. As depicted in Figure 3, in only approximately 57% of instances, Contriever identifies the oracle document as more similar to the query than the distractor documents. Figure  3 shows the distribution of entropy differences for the first token when conditioned on oracle documents versus distractor documents. It indicates that the entropy is generally lower when the response to a query is based on the oracle document rather than the distractor documents.
@@
Table 1: Performance (%) comparison of different decoding methods on benchmark datasets. "Gain" refers to the average absolute improvement (%) across all datasets and different numbers of retrieved documents when compared to the naive baseline.
Moreover, the proposed CLeHe derived from LeEns further enhances the accuracy of responses in LLMs, particularly evident in the Llama2-7B and Llama-2-13B. This indicates that contrasting low-entropy distributions of contextual knowledge with high-entropy distributions of parameterized knowledge can further strengthen LLMs’ understanding of context.
@@ 4.2 Open-Domain Question Answering
We evaluate our proposed method using four open-domain QA datasets. Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), WebQ (Berant et al., 2013) and PopQA (Mallen et al., 2023). Natural Questions includes real anonymized queries from Google’s search engine. We utilize a filtered test set (Lee et al., 2019) of 3,610 samples with answers limited to no more than five tokens. TriviaQA comprises trivia question-answer pairs that were scraped from the web. We evaluate its development set containing 7,993 samples. WebQ consists of questions generated through the Google Suggest API, with answers that are entities in Freebase. We use its test set of 2,032 samples for evaluation. PopQA is a novel entity-centric open-domain QA dataset that spans a wide range of entity popularity, emphasizing long-tail knowledge. We utilize its test set which includes 14,267 samples for evaluation. For each dataset, we retain only the questions and their corresponding answers. DPR (Karpukhin et al., 2020) is employed to retrieve the top-k passages from the Wikipedia corpus (Dec. 20, 2018) via as evidence documents for each question. Specifically, we report the performance of different decoding methods when retrieving the top-5, top-10, and top-20 documents. Exact Match (EM) metric is utilized for performance evaluation.
@@ 4.2.1 Overall Performance
4.2.2 Comparative Analysis with CAD
Our proposed CLeHe shares partial similarities with CAD, notably both employing contrastive decoding to suppress parametric knowledge. The CAD method denotes contrasting the distribution derived from Naive LLM-RAG when considering retrieved documents, against the last-layer distribution when the context is not considered. Table 2 shows the performance comparison between our CLeHe and the CAD. It’s evident that across most datasets and LLMs, the proposed CLeHe significantly outperforms the CAD, indicating that it is more effective at understanding context.
Model Method NQ TQA WebQ
Llama-2-7B CAD 38.92 65.92 30.77
CLeHe (Ours) 37.62 69.56 36.12
Llama-2-13B CAD 39.43 68.96 33.76
CLeHe (Ours) 37.31 70.12 38.19
Mistral-7B-v0.1 CAD 44.32 70.27 41.49
CLeHe (Ours) 46.32 78.26 49.76
Llama-3-8B CAD 51.80 79.01 46.26
CLeHe (Ours) 52.02 81.78 51.62
Table 2: Performance under the top-5 document setting.
4.2.3 Analysis of Layer Selection Strategies
Figure 4: Effect of layer selection strategies on WebQ under the top-5 document setting.
In our CLeHe, we dynamically select the layer with the highest entropy distribution (Eq.(6)) as the representative for parameterized knowledge at each generation time step. A simpler alternative method involves using the last layer to represent the parametric knowledge. A more complete strategy is based on distribution differences (Chuang et al., 2023) which has been demonstrated to enhance the factuality of LLMs. For this strategy, we first calculate the JSD (Jensen-Shannon Divergence) between the low-entropy ensemble distribution (Eq.(3)) and the distribution of each layer when without context. Then the distribution with the highest JSD value for contrast is selected for contrast. As shown in Figure 4, our proposed entropy-based layer selection strategy surpasses the fixed final-layer strategy. Also, an intriguing observation is that the performance of this entropy-based dynamic selection strategy slightly exceeds the best performance achievable by fixed static layers. Appendix A.1 visualizes the average entropy of tokens generated when using different layers for contrast, which shows that layers with higher entropy often correspond to better performance. Meanwhile, it performs better overall than the dynamic JSD-based strategy. It’s noteworthy that in our setting, using the JSD-based selection strategy for contrast is unstable and results in considerable performance degradation in Llama-2-13B.
@@ 4.2.3 Analysis of Layer Selection Strategies
4.2.4 Hyper-Parameter Analysis
@@ 6 Limitations
One limitation of our study is that we only validated the effectiveness of our method on question answering datasets, without testing it on other knowledge-intensive tasks such as fact verification. Extending the method proposed in this paper to other retrieval-augmented scenarios will be a future research direction.
Additionally, due to computational power constraints, we only tested the effectiveness of the proposed method on models with fewer than 13B parameters. However, whether the method proposed in this paper is applicable to LLMs with more parameters (e.g., 70B or more) remains to be explored in future research.

### v2

@@ Abstract
Augmenting Large Language Models (LLMs) with retrieved external knowledge has proven effective in improving the factual accuracy of generated responses. Despite their success, retrieval-augmented LLMs still face the distractibility issue, where the generated responses are negatively influenced by noise from both external and internal knowledge sources. In this paper, we introduce a novel, training-free decoding method guided by entropy considerations to mitigate this issue. Our approach utilizes entropy-based document-parallel ensemble decoding to prioritize low-entropy distributions from retrieved documents, thereby enhancing the extraction of relevant information of context. Additionally, it incorporates a contrastive decoding mechanism that contrasts the obtained low-entropy ensemble distribution with the high-entropy distribution derived from the model’s internal knowledge across layers, which ensures a greater emphasis on reliable external information. Extensive experiments on open-domain question answering datasets demonstrate the superiority of our method. [^1] Our code is available at https://github.com/zexuanqiu/entropy-based-decoding.
@@
Zexuan Qiu1thanks: Correspondence to: zxqiu22@cse.cuhk.edu.hk.  Zijing Ou2 Bin Wu3 Jingjing Li1 Aiwei Liu4 Irwin King1 1The Chinese University of Hong Kong 2Imperial College London 3University College London 4Tsinghua University
@@ 1 Introduction
How to eliminate the impact of the above-mentioned distractibility issue, so as to extract useful knowledge from the retrieved context for the input query, is our research focus. Although existing works strive to effectively leverage the retrieved context by directly fine-tuning retrieval-augmented LLMs (Lin et al., 2023; Zhang et al., 2024; Li et al., 2024) or incorporating trainable encoder modules (Izacard and Grave, 2020; Yen et al., 2024), these approaches require additional training, rendering them potentially impractical in resource-constrained environments. In this paper, we propose a novel decoding method guided by entropy considerations to simultaneously mitigate the impact of noisy information from both the external context and parametric knowledge. The proposed method can be seamlessly integrated into LLMs without requiring additional tuning.
@@ 2.1 Entropy-Based Document Ensemble
There are multiple choices to compute the weights. A straightforward option is to use uniform weighting, i.e., $w_{j,t}=\frac{1}{|D|}$. However, this method may fail to effectively extract valuable information when irrelevant documents are included among the top-$K$ retrieved documents. Another option is to utilize a similarity score $s(d_{j},x)$ between the query and the retrieved document, e.g., BM25, as the time-independent ensemble weights $w_{j,t}\propto s(d_{j},x)$ (Lewis et al., 2020). However, this retriever-based scoring approach may hinder the LLM’s ability to extract relevant information when the retrieved documents contain numerous distractors, drastically reducing the factual accuracy of the responses as illustrated in § 4.1.
@@ 3 Related Works
The idea of contrastive decoding (CD) has been previously applied in controllable text generation to produce non-toxic by DExperts (Liu et al., 2021). Later, Li et al. (2022) formalized CD as a method to enhance open-ended text generation without any additional training by maximizing the difference in log probabilities between an expert LLM and an amateur LLM. This approach has demonstrated strong performance in various domains, including reasoning (O’Brien and Lewis, 2023) and neural machine translation (Waldendorf et al., 2024). CD can also be interpreted as maximizing pointwise mutual information (PMI), which has proven effective in other scenarios. For instance, Li et al. (2016) uses a training objective that maximizes PMI to generate more diverse conversational responses, while CAD (Shi et al., 2023b) employs a PMI-adjusting distribution to resolve the knowledge conflict. Chuang et al. (2023) proposes a decoding strategy that contrasts different layers of the same LLM to more effectively highlight factual knowledge. Similar principles are also applied in visual LLMs, where Leng et al. (2023) mitigates object hallucination by contrasting distributions derived from original and distorted visual inputs. Alternatively, our proposed CLeHe leverages layer-wise entropy-based contrastive decoding to prioritize external knowledge over the parametric knowledge inherent in the LLM itself.
@@ 4 Experiments
We mainly compare three training-free baselines, two of which are ensemble-based methods. Naive: this method concatenates all retrieved documents directly along with the question to form the prompt for LLMs; RePlug (Shi et al., 2023c): it utilizes a normalized retriever weight to ensemble during the decoding process; AvgEns : it follows the formulation in Eq.(2) and assigns the same weight to each document during each generation step. In § 4.2.2, we also incorporate two contrastive decoding methods (i.e., CAD (Shi et al., 2023b) and DoLa (Chuang et al., 2023)) into our setting and compare with them to validate the effectiveness of our method. Experiments are conducted on 4 LLMs: Llama-2-7B, Llama-2-13B, Mistral-7B-v0.1, and Llama3-8B. Among them, Llama-3-8B supports an 8K context length, while the other three LLMs support a 4K context length. Few evaluation samples exceed 4k in context length, and they are removed to ensure the validity of the assessment.
@@
Our method introduces two hyperparameters: $\tau$ to control the relative importance of different documents during decoding; and $\beta$ balancing contextual and parametric knowledge. We extract a subset from the WebQ training set for validation to determine the hyperparameter value for each LLM. Ultimately, $\tau$ is set as 0.25 for Llama-3-8B and as 0.1 for the other three models. For $\beta$, 5.0 is chosen for Llama-2-7B, while the other models are assigned a value of 0.25. During the inference, greedy decoding is utilized for reproducibility. When looking for the layer with the highest entropy, we focus our search exclusively on the candidate layers. In our preliminary experiments, this approach improves computational efficiency and slightly enhances model performance. For Llama-2-7B, Mistral-7B-v0.1, and Llama-3-8B with 32 hidden layers, the candidate layers are set to $\{17,\dots,32\}$, and only even-numbered layers will be searched. For Llama-2-13B with 40 hidden layers, the candidate layers are set to $\{31,\dots,40\}$. All experiments are conducted on a single A100 80GB GPU.
@@ 4.1 Analyzing the Distraction Phenomenon in Retrieved Context
We are particularly interested in a challenging QA scenario proposed by  (Liu et al., 2024), in which the oracle document is surrounded by numerous semantically similar distractor documents. This configuration challenges the efficacy of the Naive LLM-RAG, preventing it from accurately deriving answers based on contextual information, thereby leading to the “Lost in the middle” phenomenon. Following (Liu et al., 2024), given a query from NaturalQuestions-Open (Kwiatkowski et al., 2019), we select a Wikipedia paragraph containing the answer from the NaturalQuestions annotations as the oracle document. Then, Contriever (Gautier et al., 2022) is employed to extract K-1 additional paragraphs from the Wikipedia corpus that are highly relevant to the query yet do not include the ground truth answer, functioning as distractor documents. The query, the oracle document, and K-1 distractor documents are subsequently processed by the LLM to generate an answer.
From Figure 2, it is evident that the performance of the Naive method which concatenates all documents to compose the context, is highly sensitive to the placement of the oracle document within the retrieved documents. Among the four evaluated LLMs, the performance of the Naive method significantly deteriorates when the oracle document is neither at the very beginning nor at the end. In contrast, since the proposed LeEns  processes each document in parallel during decoding, its performance is naturally independent of the position of the oracle document. In almost all positions, LeEns substantially surpasses the performance of the Naive method. Notably, in this challenging scenario, RePlug which ensemble documents’ distributions based on retriever weights perform exceedingly poorly, achieving results merely on par with AvgEns . Based on this observation, we further conduct the weight analysis in Figure 3. As depicted in Figure 3, in only approximately 57% of instances, Contriever identifies the oracle document as more similar to the query than the distractor documents. Figure  3 shows the distribution of entropy differences for the first token when conditioned on oracle documents versus distractor documents. It indicates that the entropy is generally lower when the response to a query is based on the oracle document rather than the distractor documents. Moreover, the proposed CLeHe derived from LeEns further enhances the accuracy of responses in LLMs, particularly evident in the Llama2-7B and Llama-2-13B. This indicates that contrasting low-entropy distributions of contextual knowledge with high-entropy distributions of parameterized knowledge can further strengthen LLMs’ understanding of context.
@@
Table 1: Performance (%) comparison of different ensemble-based methods on benchmark datasets. "Gain" refers to the average absolute improvement (%) across all datasets and different numbers of retrieved documents when compared to the naive baseline.
@@ 4.2 Open-Domain Question Answering
We evaluate our proposed method using four open-domain QA datasets. Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), WebQ (Berant et al., 2013) and PopQA (Mallen et al., 2023). Natural Questions includes real anonymized queries from Google’s search engine. We utilize a filtered test set (Lee et al., 2019) of 3,610 samples with answers limited to no more than five tokens. TriviaQA comprises trivia question-answer pairs that were scraped from the web. We evaluate its development set containing 7,993 samples. WebQ consists of questions generated through the Google Suggest API, with answers that are entities in Freebase. We use its test set of 2,032 samples for evaluation. PopQA is a novel entity-centric open-domain QA dataset that spans a wide range of entity popularity, emphasizing long-tail knowledge. We utilize its test set which includes 14,267 samples for evaluation. For each dataset, we retain only the questions and their corresponding answers. DPR (Karpukhin et al., 2020) is employed to retrieve the top-k passages from the Wikipedia corpus (Dec. 20, 2018) via as evidence documents for each question. Specifically, we report the performance of different decoding methods when retrieving the top-5, top-10, and top-20 documents. Following (Liu et al., 2024), exact match accuracy is utilized for performance evaluation.
@@ 4.2.1 Overall Performance
4.2.2 Ablation Study
Llama-2-7B Llama-2-13B Mistral-7B-v0.1 Llama-3-8B
NQ TQA WebQ NQ TQA WebQ NQ TQA WebQ NQ TQA WebQ
Naive 19.56 52.53 16.04 37.98 66.82 29.83 46.20 73.53 47.69 52.35 79.92 50.49
 w/ JSD (DoLa) 38.72 68.15 35.38 27.34 45.15 21.90 46.20 74.42 46.80 52.30 81.36 49.54
 w/ Last_Layer (CAD) 38.92 65.92 30.77 41.52 68.96 33.76 44.32 70.27 41.49 51.80 79.01 46.26
 w/ Entropy 41.36 70.32 37.30 39.91 67.38 32.73 46.30 74.59 47.05 52.28 81.60 49.61
LeEns 25.48 61.87 20.28 36.54 71.87 36.07 46.40 78.26 49.21 51.74 81.80 52.17
 w/ JSD (DoLa) 35.84 67.11 33.75 17.22 35.17 11.02 46.29 78.19 49.16 52.04 81.73 51.82
 w/ Last_Layer 30.74 62.85 23.08 38.19 71.64 38.09 45.57 75.54 48.12 52.60 81.36 52.01
 w/ Entropy (CLeHe) 37.62 69.56 36.12 37.31 72.24 38.19 46.32 78.26 49.76 52.02 81.78 51.62
Table 2: Performance on combining different external and parametric knowledge modeling methods. Experiments are conducted under the top-5 document setting.
Within the contrastive decoding framework, we investigate the compositional effects on performances by combining different modeling techniques for external and parametric knowledge. To extract knowledge from retrieved external documents, we explore two modeling approaches.: Naive RAG and our entropy-based document ensemble modeling (LeEns). Additionally, we explore three layer-based strategies to derive parametric knowledge: (i) Last-Layer strategy. It defines parametric knowledge using the distribution from the last layer of LLMs when without retrieved context. CAD (Shi et al., 2023b) utilizes this strategy, i.e., contrasting the distribution derived from Naive RAG against the last-layer context-free distribution. (ii) JSD-based strategy. It first calculates the Jensen-Shannon Divergence (JSD) between the RAG-derived distribution and the distribution of each layer when without retrieved documents, then selects the layer with the highest JSD for contrast. (iii) Our proposed entropy-based strategy. It directly selects the layer with the highest entropy as the proxy of intrinsic knowledge. As shown in Table 2, compared to the other two layer selection strategies, the proposed entropy-based strategy consistently and significantly enhances model performance in both external knowledge modeling ways of  Naive RAG and our LeEns. This improvement is particularly notable in  Llama-2-7B. The last-layer strategy markedly impairs the performance of some LLMs such as Mistral-7B-v0.1. Moreover, in our setting, the JSD-based strategy for contrast is found unstable and results in severe performance degradation in Llama-2-13B. Appendix A.2 displays the average entropy of tokens from various layers, indicating that higher entropy layers often yield better performance. Additionally, Appendix A.1 details the hyper-parameter needed to replicate Table 2.
4.2.3 Hyper-Parameter and Latency Analysis
@@
# of Docs 5 10 20
Naive 27.51 ($\times$1.00) 29.49 ($\times$1.00) 32.56 ($\times$1.00)
LeEns 30.43 ($\times$1.11) 32.90 ($\times$1.12) 37.76 ($\times$1.16)
CLeHe 31.88 ($\times$1.16) 34.10 ($\times$1.16) 38.49 ($\times$1.18)
Table 3: Decoding latency (ms/token) of Llama-2-7B based on the number of retrieved documents as context.
As for decoding latency, Table 3 shows that compared to the Naive method, our LeEns and CLeHE increase the decoding time by factors of less than 1.18, indicating that they can be applied at a reasonable cost. Appendix A.5 shows the latency of Llama-2-13B that exhibits a similar cost trend.
@@ 6 Limitations
One limitation of our study is that we only validated the effectiveness of our method on question answering datasets, without testing it on other knowledge-intensive tasks such as fact verification. Extending the method proposed in this paper to other retrieval-augmented scenarios will be a future research direction. Additionally, due to computational power constraints, we only tested the effectiveness of the proposed method on models with fewer than 13B parameters. However, whether the method proposed in this paper is applicable to LLMs with more parameters (e.g., 70B or more) remains to be explored in future research.
Acknowledgements
The work described in this paper was partially supported by Laboratory for AI-Powered Financial Technologies, InnoHK initiative and The Government of the HKSAR. The work described in this paper was also partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23).